OpenStudy (anonymous):

A solid gold cube with edge of 2" is melted down and recast as a cone with a 3" radius base. Find the height of the cone.

hero (hero):

In general, if one object is re-shaped into another object, both objects will have the same volume. If you find the volume of the cube, you will have the volume of the cone. Use the formula for Volume of a Cone to find its height.

OpenStudy (anonymous):

lol i juss need the answer no disrespect ive been workinq all day

OpenStudy (lalaly):

\[\text{Volume of cube} = 2^3= 8\] \[\text{volume of cone} =\frac{1}{3} \pi r^2 h\] \[\frac{1}{3} \pi(3^2)(h)=8\]\[h=\frac{8 \times 3}{9 \pi}\]

hero (hero):

For the first step, I would have put Volume of Cube = Volume Of Cone

OpenStudy (anonymous):

so its 9pi

OpenStudy (lalaly):

thaats what i did mr hero

OpenStudy (lalaly):

no its \[\frac{8}{3 \pi}\]

hero (hero):

Yes, Lana, but not explicitly Johny Boy, I don't know how you got 9 pi out of that.

hero (hero):

9 pi was only part of the fraction lana found for height

OpenStudy (lalaly):

oh i;m really sorryyyyyyyyy HEro

OpenStudy (anonymous):

uqh ma braINS mushed smh i been workinq all dayyy

OpenStudy (anonymous):

so my answer iss

OpenStudy (anonymous):

Hi Everyone LOL :D

hero (hero):

Hey, it's RLD! :D

OpenStudy (lalaly):

on ur calculator :)

OpenStudy (anonymous):

LOL